OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

\[(9+2\sqrt{3})(9-2\sqrt{3})\]

OpenStudy (anonymous):

81-12

OpenStudy (anonymous):

(a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

@ksrakhilesh do you mean

OpenStudy (anonymous):

\[81-4\sqrt{3}\]

OpenStudy (anonymous):

9^2-(2root3)^2

OpenStudy (anonymous):